# Vector Quantized-Elites: Unsupervised and Problem-Agnostic Quality-Diversity Optimization

### v1

@@ VI-B Robotic Arm Results
For the robotic arm experiments, we require the robot to achieve diverse poses, filling the whole operation space. In Fig. 8, we provide a visualization of some diverse poses our algorithm achieves and the final result, where the operation space is filled. For this task, we investigate two different aspects, firstly, how the structured grid archive VQ-Elites uses compares to the unstructured archive of AURORA. Secondly, if the robotic arm system has some constraints, e.g. due to motor malfunction, and thus some joints are not fully operational, can our algorithm still find optimal solutions, and how does it compare to the classic MAP-Elites algorithm?
@@
For the first case, we find that the coverage of VQ-Elites and MAP-Elites is identical while the AURORA performs slightly worse, yet, we can see a significant difference in terms of QD score, without a trend of improvement (Fig. 6). This fact indicates that the way an unstructured archive is updated is highly likely to discard high-quality and diverse individuals in order to satisfy the minimum distance between individuals rule (answers to Question 1).
@@
The second case resembles the mobile robot in the L-shaped environment. However, in this task, the constraint is on the robot itself. Despite that, our algorithm has a completely agnostic nature due to the unsupervised learning model, and hence, it can outperform the MAP-Elites algorithm (Fig. 7). We should note that this is highly predictable because in order for the MAP-Elites to achieve a competitive performance it requires knowledge of the constraints to construct a proper behavioral descriptor grid. Lastly, AURORA performs similarly to the MAP-Elites algorithm despite its unsupervised nature, while the QD score is the worst of the three due to the way the archive is constructed, as noted previously (answers to Question 2).
@@
Figure 6: Results for the robotic arm pose reaching task for VQ-Elites, MAP-Elites, and AURORA. Each graph line represents the median and the 25th and 75th quantile over ten runs for the corresponding algorithm.
@@
Figure 7: Results for the robotic arm pose reaching task with system constraints (i.e. limitation in some joints) for VQ-Elites, MAP-Elites with hard-coded BDs, and AURORA. Each graph line represents the median and the 25th and 75th quantile over ten runs for the corresponding algorithm.
Figure 8: Top: Illustration of a subset of diverse poses (individuals) achieved from the VQ-Elites algorithm. Botton: The final VQ-Elites archive that fills the behavior space with diverse poses.
VI-C Ablation Study
@@
In order to identify whether VQ-Elites is robust to hyper-parameters we perform some ablation studies. In particular, we want to investigate how the combination of the model’s update frequency, the training epochs, and the use of the cooperation component, affect the performance of our algorithm. For these experiments, we used some extreme values for the mentioned hyperparameters to investigate our algorithm’s robustness. We perform the experiments in the L-shaped mobile robot environment. The results showcase that when we update the model after every iteration of the evolution phase and train the model for 1 epoch, the performance is suboptimal despite the use of the cooperation component (Fig. 9). This is anticipated since the behavioral space grid changes continuously, but also the model fails to converge fast enough due to the low number of gradient updates. Conversely, when we update the model every 100 iterations for 100 epochs, the performance is similar to our original hyperparameters (answers to Question 3). Although one could argue that this is more selection of hyperparameters compared to the ones we use in our main experiments, we should emphasize that updating the model after every 100 iterations raises memory and wall-time issues. For every iteration, one should store all the raw data produced by the populations, which can easily lead to reaching that maximum memory capacity, especially when working with high-dimensional data like images, while slowing down significantly the model’s update.
Furthermore, we investigate the importance of the cooperation component on our algorithms as well as on the AURORA algorithm. As stated, VQ-Elites is quite robust that the cooperation does not have any impact on its performance, since the VQ-VAE model, eventually, coverages to an optimal grid structure (answers to Question 4). On the other hand, the AURORA algorithm performs poorly without this component (Fig. 10).
@@
Figure 9: Results of the ablation study regarding the model’s update frequency, the number of training epochs, and the use of cooperation component on the mobile robot maze experiment. Each graph line represents the median and the 25th and 75th quantile over ten runs for the corresponding algorithm.
@@
VII Conclusion and Discussion
In this work, we presented the Vector Quantized-Elites, VQ-Elites, algorithm, an unsupervised and problem-agnostic variation of the MAP-Elites algorithm. Our algorithm requires no prior knowledge of the task at hand to provide high-quality and diverse solutions. Additionally, we proposed two key components, the behavioral space bounding and cooperation components, which can be integrated into other unsupervised QD algorithms to mitigate the difficulties between the evolution and the update of the unsupervised model. Our work is deployed on two robotic systems, a mobile robot, and a robotic arm, while its performance is evaluated on numerous experiments. In all scenarios, our algorithm performs as well as or better than the AURORA algorithm (even when coupled with our two novel components, AURORA), making VQ-Elites algorithm a strong candidate to tackle many optimization problems.
Nevertheless, our work does not come without its limitations. It is important to note that our approach can become very costly in terms of computation time or exhibit degraded performance. The main cause for these problems originates from the selection of the raw data and the unsupervised model’s architecture. For example, utilizing very inefficient architectures for the VQ-VAE or using memory-intensive raw data (e.g. videos) to extract the BDs as well as train the model can only hinder our approach’s efficiency. Additionally, given the raw data, the selection of the model’s architecture should be tuned since a non-optimal model will not be able to construct a meaningful latent space, and consequently, a high-quality behavioral grid, resulting in poor solutions for the task at hand.
@@ VII Conclusion and Discussion
In the future, we want to incorporate generative modeling techniques such as diffusion, or flow-matching, in order to model the distribution of solutions and generate more high-quality and diverse solutions. An ideal outcome would be to use VQ-Elites for finding the minimum amount of solutions, and then via the generative modeling techniques, we would able to fill the archive successfully without being limited by the run times of the evolutionary algorithm and its sequential nature. Lastly, another interesting direction would be an epistemic uncertainty-guided evolution frame. The intuition here lies in the fact that epistemic uncertainty decreases when the number of data we train our model increases, hence, we can use the epistemic uncertainty of the model as a way to guide the QD evolution into more uncertain behaviors, with the goal of reducing the epistemic uncertainty of the VQ-VAE model, and consequently creating the most optimal behavioral space possible.

### v2

@@ VI-B Robotic Arm Results
For the robotic arm experiments, we require the robot to achieve diverse poses, filling the whole operation space. In Fig. 9, we provide a visualization of some diverse poses our algorithm achieves. For this task, we investigate two different aspects, firstly, how the structured grid archive VQ-Elites uses compares to the unstructured archive of AURORA. Secondly, if the robotic arm system has some constraints, e.g. due to motor malfunction, and thus some joints are not fully operational, can our algorithm still find optimal solutions, and how does it compare to the classic MAP-Elites algorithm?
@@
For the first scenario, it is evident that MAP-Elites is the best of the batch in both coverage and QD score. Then VQ-Elites and AURORA follow, while VQ-Elites is slightly better. This fact indicates that the way an unstructured archive is updated is highly likely to discard high-quality and diverse individuals in order to satisfy the minimum distance between individuals rule (answers to Question 1). Out of the four algorithms, vanilla AURORA performs the worst.
@@
The second case resembles the mobile robot in the L-shaped environment. However, in this task, the constraint is on the robot itself, by reducing some joint limits. Therefore, we change the limits of joints 1 and 2 (see Fig. 9) from [2.967,2.967][-2.967,2.967] to [0.5,0.5][-0.5,0.5]. Despite that, our algorithm has a completely agnostic nature due to the unsupervised learning model, and hence, it outperforms all the other approaches by a large margin in both coverage and QD score (Fig. 8). We should note that this is highly predictable because in order for the MAP-Elites to achieve a competitive performance, it requires knowledge of the constraints to construct a proper behavior descriptor grid. Lastly, both AURORA and vanilla AURORA perform similarly to the MAP-Elites algorithm despite its unsupervised nature, while the QD score of the vanilla variation is the worst of the three due to the way the archive is constructed, as noted previously (answers to Question 2). Focusing on the effective diversity and coverage diversity score, our algorithm is still outperforming the others. For instance, the $80\%$ of VQ-Elites’s archive is filled with unique behaviors in the former metric while achieving a near-perfect score in the latter. Surprisingly, the AURORA and vanilla AURORA fail to provide a competent amount of diverse behaviors.
@@
Figure 7: Results for the robotic arm pose reaching task for VQ-Elites, MAP-Elites, AURORA, and AURORA. Each graph line represents the median and the 25th and 75th quantiles over ten runs for the corresponding algorithm.
@@
Figure 8: Results for the robotic arm pose reaching task with system constraints (i.e. limitation in some joints) for VQ-Elites, MAP-Elites with hard-coded BDs, and AURORA. Each graph line represents the median and the 25th and 75th quantiles over ten runs for the corresponding algorithm.
Figure 9: Illustration of a subset of diverse poses (individuals) achieved from the VQ-Elites algorithm. The blue sphere indicates the end-effector target. Left: Robot’s starting pose with indicators about each joint’s position. Right: Learned behaviors.
VI-C MiniGrid Results
In the MiniGrid experiment, we evaluate the performance of the algorithms in a discrete and challenging exploration setting. The results, shown in Fig. 11, reveal that convergence in terms of coverage is significantly slower than in previous tasks. Notably, while VQ-Elites converges more slowly than the other methods, it achieves near-perfect effective diversity and coverage diversity (answers to Question 5 and 6). In contrast, both variants of the AURORA algorithm converge faster but fail to produce a truly diverse archive.
This difference in performance is primarily due to the use of the unstructured archive in AURORA. Such archives can be advantageous in difficult exploration scenarios when coverage alone is the primary objective. Their lack of structure allows a large number of individuals to persist in the population, effectively ”brute-forcing” the evolutionary search. This increases the pool of candidates available for crossover and mutation, which in turn improves the likelihood of producing individuals capable of better traversing the environment. This effect is particularly pronounced in the MiniGrid setting, where the high-dimensional policy space requires significantly more exploration than in the previous tasks. As a result, unstructured methods can gain an advantage in raw coverage at the cost of behavioral diversity.
Figure 10: Mean Grid Fitness of the MiniGrid exploration task for VQ-Elites, MAP-Elites with hard-coded BDs, and AURORA
Figure 11: Results for the MiniGrid exploration task for VQ-Elites, MAP-Elites with hard-coded BDs, and AURORA. Each graph line represents the median and the 25th and 75th quantiles over five runs for the corresponding algorithm.
VI-D Ablation Study
@@
In order to identify whether VQ-Elites is robust to hyper-parameters we perform some ablation studies. In particular, we want to investigate how the combination of the model’s update frequency, the training epochs, and the use of the cooperation component, affect the performance of our algorithm. For these experiments, we used some extreme values for the mentioned hyperparameters to investigate our algorithm’s robustness. We perform the experiments in the L-shaped mobile robot environment. The results showcase that when we update the model after every iteration of the evolution phase and train the model for 1 epoch, the performance is suboptimal despite the use of the cooperation component (Fig. 12). This is anticipated since the behavior space grid changes continuously, but also the model fails to converge fast enough due to the low number of gradient updates. Conversely, when we update the model every 100 iterations for 100 epochs, the performance is similar to our original hyperparameters (answers to Question 3). Although one could argue that this is more selection of hyperparameters compared to the ones we use in our main experiments, we should emphasize that updating the model after every 100 iterations raises memory and wall-time issues. For every iteration, one should store all the raw data produced by the populations, which can easily lead to reaching that maximum memory capacity, especially when working with high-dimensional data like images, while significantly slowing down the model’s update.
@@
Figure 12: Results of the ablation study regarding the model’s update frequency, the number of training epochs, and the use of the cooperation component on the mobile robot maze experiment. Each graph line represents the median and the 25th and 75th quantiles over ten runs for the corresponding algorithm.
Furthermore, we investigate the importance of the cooperation component on our algorithms as well as on the AURORA algorithm. As stated, VQ-Elites is quite robust that the cooperation does not have any impact on its performance at convergence, since the VQ-VAE model, eventually, converges to an optimal grid structure (answers to Question 4). As expected, while the cooperation mechanism is active, the QD score may fluctuate or remain low, since the algorithm intentionally allows lower-quality solutions to enter the archive in favor of promoting diversity. On the other hand, the AURORA algorithm performs poorly without this component (Fig. 13).
@@
Lastly, we performed an ablation study on the importance of behavior space bound in our framework, and through that, a very interesting result emerged. Essentially, it is evident from the Fig. 14 that without the behavior bounding our approach’s performance is hindered significantly. However, we investigated another option, which is to remove behavior space bounding but initialize the codebook of the VQ-VAE to [1,1][-1,1]. By doing so, we are able to achieve similar performance with our proposed algorithm. This indicates that providing a structured codebook to the model boosts the convergence of the model substantially, since the centers are well dispersed leading to much stronger commitment loss even in the initial stages of the learning.
Figure 14: Results of the ablation study regarding the behavior space bounding component. Each graph line represents the median and the 25th and 75th quantiles over ten runs for the corresponding algorithm.
VII Discussion and Future Work
While VQ-Elites demonstrates promising performance across domains, several open questions and limitations merit further reflection. One such issue is the observed degradation in QD performance during phases where the cooperation mechanism is active. Although cooperation facilitates faster exploration and enables model improvement through diverse samples, it can temporarily lower the archive’s overall quality, as lower-performing individuals are admitted to encourage diversity. Balancing this trade-off between diversity promotion and performance degradation remains an open challenge. Nevertheless, VQ-Elites in general does not need the cooperation mechanism, and thus does not suffer from the above limitation.
Another limitation concerns the scalability of the approach to very high-dimensional behavior spaces, especially when the raw descriptors are unstructured. While VQ-Elites effectively compresses these spaces using a learned representation, the VQ-VAE model may struggle to disentangle meaningful features in complex environments without careful design. This limitation could potentially be mitigated by incorporating explicit structure into the latent space. For instance, enforcing Riemannian [37] or $SO(3)$ manifold constraints may lead to more meaningful and interpretable representations for behaviors with geometric or rotational properties. Such structure could be introduced either through architectural constraints or auxiliary objective functions during training.
Computational requirements are another important consideration. The combined cost of online training, archive updates, and behavior space encoding can be substantial, especially for real-time or hardware-in-the-loop applications. Although VQ-Elites is compatible with asynchronous evaluation and model updates, future work should explore strategies for reducing model complexity, leveraging distillation, or introducing sparse model updates.
Moreover, while VQ-Elites was designed to be general and task-agnostic, this generality may come at the cost of suboptimality in highly structured domains. Future research could explore hybrid designs where partial task knowledge or hand-crafted descriptors are used to guide the model in early phases, later transitioning to fully learned representations as more data becomes available. Additionally, techniques such as attention mechanisms [38] or multi-modal inputs (e.g., fusing proprioceptive and visual data) could enhance representation quality in challenging domains.
Our results also reveal broader challenges in evaluating and designing unsupervised QD algorithms. Existing metrics, such as coverage and QD-score, provide limited insight into how well the algorithm captures the structure of the unknown behavior space. To address this, we introduced the Effective Diversity Ratio (EDR) and the Coverage Diversity Score (CDS), which offer a more nuanced perspective on solution redundancy and behavioral expressiveness. However, these metrics are still proxies, and future work is needed to develop principled evaluation tools that align with the goals of open-ended exploration and meaningful behavior diversity.
@@ VII Discussion and Future Work
Finally, in the future, we want to incorporate generative modeling techniques such as diffusion [39, 40], or flow-matching [41], in order to model the distribution of solutions and generate more high-quality and diverse solutions. An ideal outcome would be to use VQ-Elites for finding the minimum amount of solutions, and then via the generative modeling techniques, we would able to fill the archive successfully without being limited by the run times of the evolutionary algorithm and its sequential nature. Lastly, another interesting direction would be an epistemic uncertainty-guided evolution frame. The intuition here lies in the fact that epistemic uncertainty decreases when the number of data we train our model increases [42], hence, we can use the epistemic uncertainty of the model as a way to guide the QD evolution into more uncertain behaviors, with the goal of reducing the epistemic uncertainty of the VQ-VAE model, and consequently creating the most optimal behavioral space possible. For this purpose, we can borrow ideas from surrogate-assisted QD algorithms [43].
VIII Conclusion
In this work, we presented the Vector Quantized-Elites (VQ-Elites) algorithm, an unsupervised and problem-agnostic variation of the MAP-Elites algorithm. Our algorithm requires no prior knowledge of the task at hand to provide high-quality and diverse solutions. Additionally, we proposed two key components, the behavioral space bounding and cooperation components, which can be integrated into other unsupervised QD algorithms to mitigate the difficulties between the evolution and the update of the unsupervised model. We also introduced two novel metrics, the Effective Diversity Ratio and the Coverage Diversity Score in order to facilitate the evaluation of unsupervised QD algrithms. Our work is deployed on a mobile robot, a robotic arm, and a MiniGrid-type game task, while its performance is evaluated on numerous experiments. In all scenarios, our algorithm performs as well as or better than the AURORA algorithm (even when coupled with our two novel components, AURORA) in terms of coverage and QD score, yet it outperforms them in the EDR and CDS categories, making it a strong candidate to tackle many optimization problems.